Observations are taken on sales of a certain mountain bike in 22 sporting goods stores. The regression model was Y total sales (thousands of dollars), X1- display floor space (square meters), X2- competitors advertising expenditures (thousands of dollars), X3 advertised price (dollars per unit). (a) Fill in the values in the table given here. (Negative values should be indicated by a minus sign. Leave no cells blank be certain to enter O wherever required. Round your t-values to 3 decimal places and p-values to 4 decimal places.) Coefficient SE calc p-value Intercept FloorSpace 1,285.3 357.1 11.322 1.57 Competing Ads 6.682 3.904 0.141230.08485 Price (b-1) What is the critical value of Students t in Appendix D for a two-tailed test at a 0.01 (Round your answer to 3 decimal places.) t-value-
Answer #1

a)

l Minitab- Untitled [Worksheet 1**] File Edit Data Calc Stat Graph Editor Tools Window Help Assistant | 日 曼 晶伯 勺c|C)會투飩14 0 @ |幻园倫@ 6 C1-T C2 C3 C4 C5 C6 Predictor Coefficient SE tcalc p-value 1 Intercept 2 FloorSpace 3 Competing Ads.68200 3.90400 -1.71158 0.102439 4 Price 1285.30000 357.10000 3.59927 0.001792 11.32200 1.57000 7.21146 0.000000 -0.14123 0.084851.66447 0.111608

b-1) The critical value of student's t for a two tailed test at \alpha = 0.01

=t_{\frac{0.01}{2},n-2}=t_{0.005,20} = 2.84534